# Nonperturbative Heavy-Flavor Transport Approach for Hot QCD Matter

### v1

@@ Abstract
Introduction. The investigation of the transport properties of hot Quantum Chromodynamics (QCD) matter as formed in ultra-relativistic heavy-ion collisions (URHICs) is at the forefront of contemporary research in nuclear physics. Key phenomenological evidence from hydrodynamic simulations of the bulk medium evolution Heinz and Snellings (2013); Shuryak (2017); Bernhard et al. (2019) and from heavy-flavor (HF) diffusion calculations Beraudo et al. (2018); Acharya et al. (2022a); He et al. (2023) indicates that pertinent transport parameters, i.e., ratio of the shear viscosity to entropy density ($\eta/s$) and of the HF diffusion coefficient to the thermal wavelength, $\mathcal{D}_{s}(2\pi T)$ ($T$: temperature), are close to lower limits conjectured from the strong-coupling limit of quantum field theory Danielewicz and Gyulassy (1985); Kovtun et al. (2005). A fundamental objective is to connect these findings to the underlying QCD forces in quark-gluon plasma (QGP). This requires the use of nonperturbative interactions in a framework that rigorously implements quantum effects as the latter are expected to be critical in the vicinity of the strong-coupling limit. To realize this objective in the context of HF spectra in URHICs requires a quantitative framework for HF transport and hadronization. Such a framework is generally believed to consist of three (or five) main components Beraudo et al. (2018): heavy-quark (HQ) diffusion in the QGP, HQ hadronization, and a reliable space-time simulation of the fireball (pre-equilibrium evolution and hadronic rescattering are expected to be less important due to a short time duration and small interaction rates, respectively).
@@
The theoretical efforts toward a comprehensive approach have made tangible progress in recent years Beraudo et al. (2018); Zhao et al. (2024) leading, e.g., to rather quantitative constraints on the HQ diffusion coefficient from combined model comparisons to the nuclear modification factor and elliptic flow of $D$-mesons Acharya et al. (2022a); Sirunyan et al. (2018). While the hadronization mechanism plays an important role in the simultaneous description of these observables, more direct constraints have been obtained by measuring additional charm hadrons, in particular $D_{s}$ mesons and $\Lambda_{c}$ baryons Acharya et al. (2023a, 2022b). Significant advances in experimental precision and breadth of HF observables are now enabling stringent tests of model descriptions, both on the phenomenological side and the underlying theory that ultimately determines the momentum and $T$-dependencies of the HQ interactions.
@@
This letter reports on recent advances by combining components from two approaches while also improving on individual components. Specifically, we synthesize microscopic HQ $T$-matrix interactions in a strongly coupled QGP (sQGP) Liu and Rapp (2018); Tang et al. (2024) plus the resonance recombination model Ravagli et al. (2009) with the Trento plus viscous-fluid dynamics approach for bulk evolution Shen et al. (2016); Bernhard et al. (2019) and HF transport that merges elastic diffusion and radiative interactions Ke et al. (2018). For the first time we employ HQ $T$-matrix interactions with an underlying potential constrained by recent lattice-QCD (lQCD) data Altenkort et al. (2024, 2024) (rather than using the internal energy He and Rapp (2020)), thereby also reproducing the equation of state from lQCD as used in the hydrodynamic evolution Bazavov et al. (2014). In addition, we refine the effective transport implementation of radiation with an improved interference behavior as found for static media Ke et al. (2019).
@@
Langevin Approach in sQGP. In the LIDO model Ke et al. (2018) the time evolution of the HQ phase space distribution, $f_{Q}(t,x,p)$, can be formally expressed as
@@
The elastic HQ transport coefficients (friction and momentum diffusion) figuring in this operator are computed from heavy-light $T$-matrices within the quantum many-body theory developed in Ref. Liu and Rapp (2018). In particular, they encode off-shell properties of the medium through the use of light-parton spectral functions that are self-consistently calculated, while the bare masses in the underlying Hamiltonian are adjusted to reproduce the lQCD equation of state within the thermodynamically conserving quantum many-body formalism of Luttinger-Ward and Baym Luttinger and Ward (1960); Baym and Kadanoff (1961). Schematically, the underlying system of Dyson-Schwinger equations can be written as
@@
For the input potential (which is universal to both heavy- and light-parton interactions and includes relativistic corrections) we use two versions that are based on two sets of constraints from lattice QCD: (a) HQ free energies with a rather strong vector component in the confining potential deduced from spin and spin-orbit splittings in the vacuum quarkonium spectra (referred to as VCP), and (b) recent Wilson line correlators (referred to as WLC) which turn out to be particularly sensitive to the collisional widths of the heavy quarks in the QGP Tang et al. (2024), but prefer a somewhat weaker momentum dependence. These features are reflected in the resulting friction coefficient, $A(p)$, displayed in the left panel of Fig. 1. The predicted diffusion coefficient, $\mathcal{D}_{s}=\frac{T}{MA(p=0)}$, shown in the right panel, agrees somewhat better with lQCD data for the VCP scenario Altenkort et al. (2023); Bollweg et al. (2025).
Figure 1: Charm-quark friction (left panel) and spatial-diffusion coefficient (right panel) from selfconsistent heavy-light $T$-matrices in a strongly coupled QGP for the VCP (red lines) and WLC (blue lines) scenarios, compared to lQCD data Altenkort et al. (2024, 2023); Bollweg et al. (2025).
Following previous studies (and supported by case studies with the Boltzmann equation Beraudo et al. (2018)), we enforce the Einstein relation by setting the momentum diffusion coefficient (also referred to as $\kappa$) from the friction coefficient via $B=TE_{p}A(p)$, where $E_{p}=(M^{2}+p^{2})^{1/2}$ is the HQ on-shell energy. The diffusing quark is treated as a quasiparticle, but the transport coefficients encode the full off-shell dynamics of the medium through the use of the broad thermal-parton spectral functions underlying the description of the equation of state (EoS) Liu et al. (2019). This is critical for accessing the interaction strength of the broad heavy-light $D$-meson bound states that form below the nominal charm-light threshold as the temperature drops toward the hadronization transition.
@@
The expression in eq. (6) applies to the Bethe-Heitler (BH) regime where the gluon formation time, $\tau_{f}$, is short compared to its mean-free-path, $\lambda$. In the Landau-Pomeranchuk-Migdal (LPM) regime where $\tau_{f}\gg\lambda$, we account for coherent scatterings with the medium. In perturbative-QCD effective kinetic theory Arnold et al. (2002, 2003), the branching rate in the deep-LPM regime is reduced by a factor $\sim$ $\lambda/\tau_{f}$ compared to the BH limit. This led the authors of Ref. Ke et al. (2019) to introduce a prescription in the transport treatment where, subsequent to a $1\to 2$ branching at time $t_{0}$, the daughter partons are not instantly considered as independent, but the parton system continues elastic scatterings with the medium until a time $t$, where
@@
The LIDO model implements the HQ dead cone effect via a suppression factor, $\left(1+\theta_{g}^{2}/\theta_{M}^{2}\right)^{-n}$, relative to the gluon radiation rate of a light quark ($\theta_{g}$: gluon emission angle, $\theta_{M}=M/E_{p}$: dead cone angle). The power exponent nn=2 was employed in earlier work Ke et al. (2019), but recent studies Fu et al. (2025) show that nn=1 provides better consistency with calculations in soft-collinear effective theory Kang et al. (2017).
@@
Heavy-quark hadronization. Due to a large ambient thermal-parton density, recombination processes play an important role in HQ hadronization at low and intermediate momentum. This has been supported early on by HF data in Au-Au(200 GeV) collisions at RHIC, especially for the elliptic flow Adare et al. (2007). The recombination of heavy quarks with thermal light quarks enables the HF hadrons to directly inherit the $v_{2}$ of the bulk medium around the hadronization transition, where it is believed to be near its maximal value. Here, we focus on the resonance recombination model (RRM) Ravagli and Rapp (2007), which possesses several effectual features. Being derived from an underlying Boltzmann equation, it conserves 4-momentum in the parton-to-hadron conversion and yields the correct equilibrium limit of the produced hadrons, also for expanding media with radial and elliptic flow He et al. (2012). Its key ingredients are resonant cross sections for heavy-light parton scattering into $D$-mesons and charmed baryons, which naturally follow from the $T$-matrix interactions used in the diffusion part described above. It thereby enables a straightforward inclusion of excited states, which are likely important for the hadro-chemistry of the HF hadrons in nuclear collisions. The RRM is supplemented by fragmentation processes which take over at high transverse momentum ($p_{\rm T}$) and in $pp$ collisions.
The RRM equation for the phase space distribution (PSD) of a meson $M$ can be obtained as Ravagli and Rapp (2007)
@@
where the resonance cross section, $\sigma(s)$, for $q+\bar{q}\to M$ is of relativistic Breit-Wigner form with a meson width $\Gamma_{M}$; $f_{q,\bar{q}}$ are the thermal-quark/antiquark PSDs, $v_{\text{rel}}$ is the relative velocity, and $\gamma_{M}=E_{M}/m_{M}$. We implement the RRM on a hydrodynamic hypersurface of the LIDO model at a hadronization temperature of $T_{H}=160$ MeV, employing an event-by-event processing of the $c$ quarks from the diffusion calculation that retains the correlations between their momentum and spatial point of hadronization He and Rapp (2020). These extend the $p_{\rm T}$ reach of the RRM component significantly and affect both open-charm He and Rapp (2020) and charmonium He et al. (2023) spectra for $p_{\rm T}\gtrsim 5$ GeV.
@@
To evaluate the partition between recombination and fragmentation, a recombination probability, $P_{\text{rec}}(p^{*}_{c})$, is determined selfconsistently from the RRM expression, eq. (9), for each $c$ quark from the diffusion simulation in the fluid rest frame with momentum $p^{*}_{c}$. After all charm-hadron states are summed up, an overall normalization is fixed to render this probability equal to one at $p_{c}^{*}$=0 (as there is no energy available for fragmentation). The masses of the light (qq=$u,d$), strange ($s$) and $c$ quarks are taken as $m_{q,s,c}=0.3,0.4,1.5$ GeV, and $\Gamma_{M}=0.1$ GeV. The RRM includes all charmed-hadron states listed by the particle data group (PDG) Tanabashi et al. (2018) plus charm-baryon states predicted by the relativistic quark model Ebert et al. (2011), with branching ratios guided by the PDG. For $c$ quarks that do not recombine, fragmentation is carried out in the lab frame with a probability $1-P_{\text{rec}}(p^{*}_{c})$ using fragmentation functions (FFs) from HQ effective theory (HQET).
@@
The FFs are determined in fits to $p_{\rm T}$ spectra of $D$, $D^{*}$, $D_{s}^{+}$ and $\Lambda_{c}$ hadrons in $pp$ collisions at the LHC (summarized in Fig. 2) following Ref. He and Rapp (2019). They are based on $c$-quark spectra from the fixed-order next-to-leading logarithm (FONLL) framework and carry hadro-chemical weights for the individual charm-hadron species determined from a statistical hadronization model with $T_{H}=170$  MeV. A single overall normalization, $N$, is applied, amounting to the total charm cross section, $d\sigma_{c\bar{c}}/dy$. We also require a strangeness suppression factor of $\gamma_{s}$=0.7 to reproduce the yield of the charm-strange $D_{s}^{+}$ mesons in $pp$ collisions. This is slightly larger than the value of 0.6 in previous work He and Rapp (2019), caused by revised branching ratios of excited $D_{s}^{*}$ whose masses are above the $DK$ threshold. Previously all these $D_{s}^{*}$’s were decayed into $D_{s}\pi$, while here, following available information from the PDG, we take a branching ratio of $\sim$80% for $D_{s}^{*}\to DK$ above the $DK$ threshold. For the RRM component in URHICs, the thermal strange-quark PSDs in the QGP are evaluated with $\gamma_{s}=1$.
@@
Figure 2: Production cross sections of prompt $D^{0}$ (red), $D_{s}^{+}$ (green) and $\Lambda_{c}^{+}$ hadrons in $pp$ collisions at $\sqrt{s_{\mathrm{NN}}}$=5.02 TeV, using FONLL $c$-quark spectra (dashed line) and HQET fragmentation with statistical weights from the SHM, compared to ALICE data Acharya et al. (2019, 2023b).
@@
Bulk Evolution Model. A realistic bulk medium evolution for URHICs is a prerequisite for a reliable simulation of HQ transport within, as it provides the local temperature and flow fields for the transport coefficients that govern HQ dynamics. In the present work, we utilize a 2+1-dimensional viscous hydrodynamical evolution, which consists of several stages. Initial conditions are generated at $\tau=0^{+}$ based on a partition of participant and binary-collision profiles Moreland et al. (2015). The pre-equilibrium stage is approximated by free streaming until the start of hydrodynamics at $\tau_{\rm fs}$ Liu et al. (2015) for the QGP evolution with an up-to-date lattice EoS Shen et al. (2016); Bazavov et al. (2014). The hydrodynamic medium is converted into an ensemble of light hadrons via sampling a hypersurface of fixed temperature of $T_{\rm CF}=152$ MeV utilizing the Cooper-Frye formula. Note that this does not have to be the same as the temperature where the medium converts from partons to hadrons, which is not sharply defined. The hadronic ensemble is subject to further rescattering and decays utilizing the Ultra-Relativistic Quantum Molecular Dynamics (UrQMD) model Bass et al. (1998); Bleicher et al. (1999). All of these stages involve parameters that have previously been calibrated to reproduce a vast array of bulk observables at the LHC Bernhard et al. (2019), providing a description of the bulk evolution of the QGP at high precision. We will, however, defer the study of the impact of both the pre-equilibrium and hadronic phase on the HF spectra Fu et al. (2025), as their effect on the nuclear modification factor and elliptic flow is expected to small Das et al. (2025); Chesler et al. (2013); Mrowczynski (2018).
@@
Figure 3: Nuclear modification factor, $R_{\rm AA}$ (lower panels), and elliptic flow, $v_{2}$ (upper panels), of $D^{0}$ mesons at mid-rapidity in Pb-Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02\text{ TeV}$, calculated from our integrated HF transport and hadronization approach using $T$-matrix interactions with VCP (red lines) or WLC (blue lines) constraints. The red band in the lower left panel illustrates a shadowing range of 70–80% (the default is 75%). Experimental $R_{\rm AA}$ and $v_{2}$ data are from Refs. Acharya et al. (2022a) and Acharya et al. (2021); Sirunyan et al. (2021), respectively.
We will focus our analysis on open HF meson production in Pb–Pb collisions at a center-of-mass energy of $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV. The time evolution of the HQ distribution function in the QGP requires initial conditions in both position and momentum space. We assume that all $c$-quarks are produced upon initial impact of the incoming Pb nuclei with a Glauber model binary-collision profile for their spatial distribution. In momentum space, we employ the FONLL Cacciari et al. (2001) spectra determined in $pp$ fits, augmented by nuclear shadowing with a $p_{\rm T}$-dependence taken from Ref. Emel’yanov et al. (1997) and an integrated suppression of 25% in semi-/central Pb-Pb collisions, compatible with recent measurements Acharya et al. (2023b) (when varying the suppression over 20-30%, the maximum in the $R_{\rm AA}$ varies by up to 5%, while the $v_{2}$ is affected very little). At the hydro thermalization time, $\tau_{\rm fs}$, we commence the QGP diffusion with elastic $T$-matrix interactions and radiation as described above, followed by the RRM-HQET hadronization model. As mentioned above, we neglect effects from the pre-equilibrium and hadronic phase in the present study.
@@
Our results for $D$-mesons are summarized in Fig. 3. Fair agreement with ALICE and CMS data at mid-rapidity is found, keeping in mind that no parameter has been tuned specifically to these data. The description is better within the VCP constraints for the $T$-matrix interactions, which is reassuring given that the underlying spatial diffusion coefficient aligns better with the pertinent lQCD data. The larger interaction strength (smaller $\mathcal{D}_{s}$) at low sQGP temperatures appears to be vital to produce sufficient $v_{2}$ around its maximum, reiterating the important role of this observable in assessing the transport properties of the QCD medium. The $v_{2}$ tends to be underestimated toward higher $p_{\rm T}$: radiative interactions become dominant but are less effective in generating $v_{2}$ compared to, e.g., the purely elastic interactions in a similar framework in Ref. He and Rapp (2020). However, as pointed out in Ref. Noronha-Hostler et al. (2016), the inclusion of event plane fluctuations, which imply a misalignment with the hydrodynamic reaction plane, may help to remedy this problem.
Figure 4: $D_{s}^{+}/D^{0}$ ratio in Pb-Pb(5.02 TeV) collisions at a hadronization temperature of $T_{H}$=160 MeV. The bands reflect a range of the strange-quark mass of $m_{s}$=0.4-0.45 GeV. ALICE data are taken from Acharya et al. (2022b, 2019)
As an example of charm-hadron chemistry in Pb-Pb collisions, we display in Fig. 4 our calculated $D_{s}^{+}/D^{0}$ ratio. An improvement over previous work He and Rapp (2020) is found, mostly originating from the revised branching ratios of excited $D_{s}^{*}$ mesons into the $DK$ channel.

### v2

@@ Abstract
Introduction. The investigation of the transport properties of hot Quantum Chromodynamics (QCD) matter as formed in ultra-relativistic heavy-ion collisions (URHICs) is at the forefront of contemporary research in nuclear physics. Key phenomenological evidence from hydrodynamic simulations of the bulk medium evolution Heinz:2013th; Shuryak:2014zxa; Bernhard:2019bmu and from heavy-flavor (HF) diffusion calculations Rapp:2018qla; ALICE:2021rxa; He:2022ywp indicates that pertinent transport parameters, i.e., ratio of the shear viscosity to entropy density ($\eta/s$) and of the HF diffusion coefficient to the thermal wavelength, $\mathcal{D}_{s}(2\pi T)$ ($T$: temperature), are close to lower limits conjectured from the strong-coupling limit of quantum field theory Danielewicz:1984ww; Kovtun:2004de. A fundamental objective is to connect these findings to the underlying QCD forces in quark-gluon plasma (QGP). This requires the use of nonperturbative interactions in a framework that rigorously implements quantum effects as the latter are expected to be critical in the vicinity of the strong-coupling limit. To realize this objective in the context of HF spectra in URHICs requires a quantitative framework for HF transport and hadronization. Such a framework is generally believed to consist of three (or five) main components Rapp:2018qla: heavy-quark (HQ) diffusion in the QGP, HQ hadronization, and a reliable space-time simulation of the fireball (pre-equilibrium evolution and hadronic rescattering are expected to be less important due to a short time duration and small interaction rates, respectively).
@@
The theoretical efforts toward a comprehensive approach have made tangible progress in recent years Rapp:2018qla; Zhao:2023nrz leading, e.g., to rather quantitative constraints on the HQ diffusion coefficient from combined model comparisons to the nuclear modification factor and elliptic flow of $D$-mesons ALICE:2021rxa; CMS:2017vhp. While the hadronization mechanism plays an important role in the simultaneous description of these observables, more direct constraints have been obtained by measuring additional charm hadrons, in particular $D_{s}$ mesons and $\Lambda_{c}$ baryons ALICE:2021bib; ALICE:2021kfc. Significant advances in experimental precision and breadth of HF observables are now enabling stringent tests of model descriptions, both on the phenomenological side and the underlying theory that ultimately determines the momentum and $T$-dependencies of the HQ interactions.
@@
This letter reports on recent advances by combining components from two approaches while also improving on individual components. Specifically, we synthesize microscopic HQ $T$-matrix interactions in a strongly coupled QGP (sQGP) Liu:2017qah; Tang:2023tkm plus the resonance recombination model Ravagli:2008rt with the Trento plus viscous-fluid dynamics approach for bulk evolution Shen:2014vra; Bernhard:2019bmu and HF transport that merges elastic diffusion and radiative interactions Ke:2018tsh. For the first time we employ HQ $T$-matrix interactions with an underlying potential constrained by recent lattice-QCD (lQCD) data Altenkort:2023eav; Altenkort:2023eav (rather than using the internal energy He:2019vgs), thereby also reproducing the equation of state from lQCD as used in the hydrodynamic evolution HotQCD:2014kol. In addition, we refine the effective transport implementation of radiation with an improved interference behavior as found for static media Ke:2018jem.
@@
Langevin Approach in sQGP. In the LIDO model Ke:2018tsh the time evolution of the HQ phase space distribution, $f_{Q}(t,x,p)$, can be formally expressed as
@@
The elastic HQ transport coefficients (friction and momentum diffusion) figuring in this operator are computed from heavy-light $T$-matrices within the quantum many-body theory developed in Ref. Liu:2017qah. In particular, they encode off-shell properties of the medium through the use of light-parton spectral functions that are self-consistently calculated, while the bare masses in the underlying Hamiltonian are adjusted to reproduce the lQCD equation of state within the thermodynamically conserving quantum many-body formalism of Luttinger-Ward and Baym Luttinger:1960ua; Baym:1961zz. Schematically, the underlying system of Dyson-Schwinger equations can be written as
@@
For the input potential (which is universal to both heavy- and light-parton interactions and includes relativistic corrections) we use two versions that are based on two sets of constraints from lattice QCD: (a) HQ free energies with a rather strong vector component in the confining potential deduced from spin and spin-orbit splittings in the vacuum quarkonium spectra (referred to as VCP), and (b) recent Wilson line correlators (referred to as WLC) which turn out to be particularly sensitive to the collisional widths of the heavy quarks in the QGP Tang:2023tkm, but prefer a somewhat weaker momentum dependence. These features are reflected in the resulting friction coefficient, $A(p)$, displayed in the left panel of Fig. 1. The predicted diffusion coefficient, $\mathcal{D}_{s}=\frac{T}{MA(p=0)}$, shown in the right panel, agrees somewhat better with lQCD data for the VCP scenario Altenkort:2023oms; HotQCD:2025fbd. The contribution from gluon radiation is negligible Liu:2020dlt.
Figure 1: Charm-quark friction (left panel) and spatial-diffusion coefficient (right panel) from selfconsistent heavy-light $T$-matrices in a strongly coupled QGP for the VCP (red lines) and WLC (blue lines) scenarios, compared to lQCD data Altenkort:2023eav; Altenkort:2023oms; HotQCD:2025fbd.
Following previous studies (and supported by case studies with the Boltzmann equation Rapp:2018qla), we enforce the Einstein relation by setting the momentum diffusion coefficient (also referred to as $\kappa$) from the friction coefficient via $B=TE_{p}A(p)$, where $E_{p}=(M^{2}+p^{2})^{1/2}$ is the HQ on-shell energy. The diffusing quark is treated as a quasiparticle, but the transport coefficients encode the full off-shell dynamics of the medium through the use of the broad thermal-parton spectral functions underlying the description of the equation of state (EoS) Liu:2018syc. This is critical for accessing the interaction strength of the broad heavy-light $D$-meson bound states that form below the nominal charm-light threshold as the temperature drops toward the hadronization transition.
@@
The expression in eq. (6) applies to the Bethe-Heitler (BH) regime where the gluon formation time, $\tau_{f}$, is short compared to its mean-free-path, $\lambda$. In the Landau-Pomeranchuk-Migdal (LPM) regime where $\tau_{f}\gg\lambda$, we account for coherent scatterings with the medium. In perturbative-QCD effective kinetic theory Arnold:2002ja; Arnold:2002zm, the branching rate in the deep-LPM regime is reduced by a factor $\sim$ $\lambda/\tau_{f}$ compared to the BH limit. This led the authors of Ref. Ke:2018jem to introduce a prescription in the transport treatment where, subsequent to a $1\to 2$ branching at time $t_{0}$, the daughter partons are not instantly considered as independent, but the parton system continues elastic scatterings with the medium until a time $t$, where
@@
The LIDO model implements the HQ dead cone effect via a suppression factor, $\left(1+\theta_{g}^{2}/\theta_{M}^{2}\right)^{-n}$, relative to the gluon radiation rate of a light quark ($\theta_{g}$: gluon emission angle, $\theta_{M}=M/E_{p}$: dead cone angle). The power exponent nn=2 was employed in earlier work Ke:2018jem, but recent studies Fu:2025 show that nn=1 provides better consistency with calculations in soft-collinear effective theory Kang:2016ofv.
@@
Heavy-quark hadronization. Due to a large ambient thermal-parton density, recombination processes play an important role in HQ hadronization at low and intermediate momentum. This has been supported early on by HF data in Au-Au(200 GeV) collisions at RHIC, especially for the elliptic flow PHENIX:2006iih. The recombination of heavy quarks with thermal light quarks enables the HF hadrons to directly inherit the $v_{2}$ of the bulk medium around the hadronization transition, where it is believed to be near its maximal value. Here, we focus on the resonance recombination model (RRM) Ravagli:2007xx, which possesses several effectual features. Being derived from an underlying Boltzmann equation, it conserves 4-momentum in the parton-to-hadron conversion and yields the correct equilibrium limit of the produced hadrons, also for expanding media with radial and elliptic flow He:2011qa. Its key ingredients are resonant cross sections for heavy-light parton scattering into $D$-mesons and charmed baryons, which naturally follow from the $T$-matrix interactions used in the diffusion part described above. It thereby enables a straightforward inclusion of excited states, which are likely important for the hadro-chemistry of the HF hadrons in nuclear collisions. The RRM is supplemented by fragmentation processes which take over at high transverse momentum ($p_{\rm T}$) and in $pp$ collisions.
The RRM equation for the phase space distribution (PSD) of a meson $M$ can be obtained as Ravagli:2007xx
@@
where the resonance cross section, $\sigma(s)$, for $q+\bar{q}\to M$ is of relativistic Breit-Wigner form with a meson width $\Gamma_{M}$; $f_{q,\bar{q}}$ are the thermal-quark/antiquark PSDs, $v_{\text{rel}}$ is the relative velocity, and $\gamma_{M}=E_{M}/m_{M}$. We implement the RRM on a hydrodynamic hypersurface of the LIDO model at a hadronization temperature of $T_{H}=160$ MeV, employing an event-by-event processing of the $c$ quarks from the diffusion calculation that retains the correlations between their momentum and spatial point of hadronization He:2019vgs. These extend the $p_{\rm T}$ reach of the RRM component significantly and affect both open-charm He:2019vgs and charmonium He:2022ywp spectra for $p_{\rm T}\gtrsim 5$ GeV.
@@
To evaluate the partition between recombination and fragmentation, a recombination probability, $P_{\text{rec}}(p^{*}_{c})$, is determined selfconsistently from the RRM expression, eq. (9), for each $c$ quark from the diffusion simulation in the fluid rest frame with momentum $p^{*}_{c}$. After all charm-hadron states are summed up, an overall normalization is fixed to render this probability equal to one at $p_{c}^{*}$=0 (as there is no energy available for fragmentation). The masses of the light (qq=$u,d$), strange ($s$) and $c$ quarks are taken as $m_{q,s,c}=0.3,0.4,1.5$ GeV, and $\Gamma_{M}=0.1$ GeV. The RRM includes all charmed-hadron states listed by the particle data group (PDG) ParticleDataGroup:2018ovx plus charm-baryon states predicted by the relativistic quark model Ebert:2011kk, with branching ratios guided by the PDG. For $c$ quarks that do not recombine, fragmentation is carried out in the lab frame with a probability $1-P_{\text{rec}}(p^{*}_{c})$ using fragmentation functions (FFs) from HQ effective theory (HQET).
@@
The FFs are determined in fits to $p_{\rm T}$ spectra of $D$, $D^{*}$, $D_{s}^{+}$ and $\Lambda_{c}$ hadrons in $pp$ collisions at the LHC (summarized in Fig. 2) following Ref. He:2019tik. They are based on $c$-quark spectra from the fixed-order next-to-leading logarithm (FONLL) framework and carry hadro-chemical weights for the individual charm-hadron species determined from a statistical hadronization model with $T_{H}=170$  MeV. A single overall normalization, $N$, is applied, amounting to the total charm cross section, $d\sigma_{c\bar{c}}/dy$. We also require a strangeness suppression factor of $\gamma_{s}$=0.7 to reproduce the yield of the charm-strange $D_{s}^{+}$ mesons in $pp$ collisions. This is slightly larger than the value of 0.6 in previous work He:2019tik, caused by revised branching ratios of excited $D_{s}^{*}$ whose masses are above the $DK$ threshold. Previously all these $D_{s}^{*}$’s were decayed into $D_{s}\pi$, while here, following available information from the PDG, we take a branching ratio of $\sim$80% for $D_{s}^{*}\to DK$ above the $DK$ threshold. For the RRM component in URHICs, the thermal strange-quark PSDs in the QGP are evaluated with $\gamma_{s}=1$.
@@
Figure 2: Production cross sections of prompt $D^{0}$ (red), $D_{s}^{+}$ (green) and $\Lambda_{c}^{+}$ (blue) hadrons in $pp$ collisions at $\sqrt{s_{\mathrm{NN}}}$=5.02 TeV, using FONLL $c$-quark spectra (dashed line) and HQET fragmentation with statistical weights from the SHM, compared to ALICE data ALICE:2019nxm; ALICE:2022exq.
@@
Bulk Evolution Model. A realistic bulk medium evolution for URHICs is a prerequisite for a reliable simulation of HQ transport within, as it provides the local temperature and flow fields for the transport coefficients that govern HQ dynamics. In the present work, we utilize a 2+1-dimensional viscous hydrodynamical evolution, which consists of several stages. Initial conditions are generated at $\tau=0^{+}$ based on a partition of participant and binary-collision profiles Moreland:2014oya. The pre-equilibrium stage is approximated by free streaming until the start of hydrodynamics at $\tau_{\rm fs}$ Liu:2015nwa for the QGP evolution with an up-to-date lattice EoS Shen:2014vra; HotQCD:2014kol. The hydrodynamic medium is converted into an ensemble of light hadrons via sampling a hypersurface of fixed temperature of $T_{\rm CF}=152$ MeV utilizing the Cooper-Frye formula. Note that this does not have to be the same as the temperature where the medium converts from partons to hadrons, which is not sharply defined. The hadronic ensemble is subject to further rescattering and decays utilizing the Ultra-Relativistic Quantum Molecular Dynamics (UrQMD) model Bass:1998ca; Bleicher:1999xi. All of these stages involve parameters that have previously been calibrated to reproduce a vast array of bulk observables at the LHC Bernhard:2019bmu, providing a description of the bulk evolution of the QGP at high precision. We will, however, defer the study of the impact of both the pre-equilibrium and hadronic phase on the HF spectra Fu:2025, as their effect on the nuclear modification factor and elliptic flow is expected to small Das:2024vac; Chesler:2013cqa; Mrowczynski:2017kso.
@@
Figure 3: Nuclear modification factor, $R_{\rm AA}$ (lower panels), and elliptic flow, $v_{2}$ (upper panels), of $D^{0}$ mesons at mid-rapidity in Pb-Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02\text{ TeV}$, calculated from our integrated HF transport and hadronization approach using $T$-matrix interactions with VCP (red lines) or WLC (blue lines) constraints. The red band in the lower left panel illustrates a shadowing range of 70–80% (the default is 75%). Experimental $R_{\rm AA}$ and $v_{2}$ data are from Refs. ALICE:2021rxa; CMS:2017qjw and ALICE:2020iug; CMS:2020bnz, respectively. The dashed lines in the left panels are for $c$-quarks just prior to hadronization.
We will focus our analysis on open HF meson production in Pb–Pb collisions at a center-of-mass energy of $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV. The time evolution of the HQ distribution function in the QGP requires initial conditions in both position and momentum space. We assume that all $c$-quarks are produced upon initial impact of the incoming Pb nuclei with a Glauber model binary-collision profile for their spatial distribution. In momentum space, we employ the FONLL Cacciari:2001td spectra determined in $pp$ fits, augmented by nuclear shadowing with a $p_{\rm T}$-dependence taken from Ref. Emelyanov:1997guf and an integrated suppression of 25% in semi-/central Pb-Pb collisions, compatible with recent measurements ALICE:2022exq (when varying the suppression over 20-30%, the maximum in the $R_{\rm AA}$ varies by up to 5%, while the $v_{2}$ is affected very little). At the hydro thermalization time, $\tau_{\rm fs}$, we commence the QGP diffusion with elastic $T$-matrix interactions and radiation as described above, followed by the RRM-HQET hadronization model. As mentioned above, we neglect effects from the pre-equilibrium and hadronic phase in the present study.
@@
Our results for $D$-mesons are summarized in Fig. 3. Fair agreement with ALICE and CMS data at mid-rapidity is found, keeping in mind that no parameter has been tuned specifically to these data. The description is better within the VCP constraints for the $T$-matrix interactions, which is reassuring given that the underlying spatial diffusion coefficient aligns better with the pertinent lQCD data. The larger interaction strength (smaller $\mathcal{D}_{s}$) at low sQGP temperatures appears to be vital to produce sufficient $v_{2}$ around its maximum, reiterating the important role of this observable in assessing the transport properties of the QCD medium. Hadronization remains a key component in the description, significantly augmenting the $v_{2}$ and reshaping the $R_{\rm AA}$ (cf. the dashed lines in the left panel of Fig. 3). Toward higher $p_{\rm T}$ the $v_{2}$ tends to be underestimated: radiative interactions become dominant but are less effective in generating $v_{2}$ compared to, e.g., the purely elastic interactions in a similar framework in Ref. He:2019vgs. However, as pointed out in Ref. Noronha-Hostler:2016eow, the inclusion of event plane fluctuations, which imply a misalignment with the hydrodynamic reaction plane, may help to remedy this problem. Thus, the combination of nonperturbative interactions at low and intermediate $p_{\rm T}$ yielding to radiation of massive gluons at high $p_{\rm T}$ appears to be a viable scenario (see, e.g., Refs. Gossiaux:2010yx; Uphoff:2014hza; Grishmanovskii:2025mnc for related discussions).
Figure 4: $D_{s}^{+}/D^{0}$ ratio in Pb-Pb(5.02 TeV) collisions at a hadronization temperature of $T_{H}$=160 MeV. The bands reflect a range of the strange-quark mass of $m_{s}$=0.4-0.45 GeV. ALICE data are taken from ALICE:2021kfc; ALICE:2019nxm
As an example of charm-hadron chemistry in Pb-Pb collisions, we display in Fig. 4 our calculated $D_{s}^{+}/D^{0}$ ratio. An improvement over previous work He:2019vgs is found, mostly originating from the revised branching ratios of excited $D_{s}^{*}$ mesons into the $DK$ channel.